JNTU

JNTU B.Tech I Year MATHEMATICS - III Previous Question Paper (August - 2018 - R16)

B.Tech | I Year | I-II | R16 | August - 2018

Answer PaperPaper Code: 132ACMaximum Marks: 75Time: 3 hours

Note: This question paper contains two parts A and B. Part A is compulsory which carries 25 marks. Answer all questions in Part A. Part B consists of 5 Units. Answer any one full question from each unit. Each question carries 10 marks and may have a, b, c as sub questions.

Part A

1a

Find the mean and variance of the probability distribution having pdf,

f(x) = e^(-x), x > 0

2 marks

Answer

For the probability density function f(x) = e^(-x) where x > 0, the mean (E[X]) is calculated as the integral from 0 to infinity of x * e^(-x) dx. Using integration by parts, this yields 1. The variance (V[X]) is calculated as E[X^2] - (E[X])^2. E[X^2] is the integral from 0 to infinity of x^2 * e^(-x) dx, which equals 2. Therefore, the variance is 2 - (1)^2 = 1.

1b

A fair coin is tossed until a head or consecutive five tails occurs. Find the discrete probability distribution.

3 marks

Answer

Let X be the random variable representing the number of tosses until either a head (H) appears or five consecutive tails (TTTTT) occur. The possible values for X are 1, 2, 3, 4, and 5.

We calculate the probability P(X = x) for each outcome:

  • P(X = 1): The first toss is H. P(X = 1) = 1/2.
  • P(X = 2): The first toss is T, second is H. P(X = 2) = (1/2) * (1/2) = 1/4.
  • P(X = 3): The sequence is TTH. P(X = 3) = (1/2)^3 = 1/8.
  • P(X = 4): The sequence is TTTH. P(X = 4) = (1/2)^4 = 1/16.
  • P(X = 5): The sequence is TTTTH or TTTTT. P(X = 5) = (1/2)^5 + (1/2)^5 = 1/32 + 1/32 = 2/32 = 1/16.

The probability distribution is given by P(X=1)=1/2, P(X=2)=1/4, P(X=3)=1/8, P(X=4)=1/16, and P(X=5)=1/16. Note that the sum of probabilities is 1/2 + 1/4 + 1/8 + 1/16 + 1/16 = 1.

1c

Write the conditions of validity of

χ²-test
.

2 marks

Answer

Sign in to view the complete answer paper.

1d

Construct sampling distribution of means for the population 3, 7, 11, 15 by drawing samples of size two without replacement. Determine i)

μ
ii)
σ
iii) Sampling distribution of means.

3 marks

Answer

Sign in to view the complete answer paper.

1e

Discuss types of errors of the test of hypothesis.

2 marks

Answer

Sign in to view the complete answer paper.

1f

Give the graphical interpretation of the bisection method.

3 marks

Answer

Sign in to view the complete answer paper.

1g

Write the iterative formula for finding

N^(1/3)
using method of false position.

2 marks

Answer

Sign in to view the complete answer paper.

1h

Explain briefly about method of least square.

3 marks

Answer

Sign in to view the complete answer paper.

1i

Derive Trapezoidal rule for computing integral

∫f(x)dx
from a to b.

2 marks

Answer

Sign in to view the complete answer paper.

1j

What are the limitations of Taylor’s series method?

3 marks

Answer

Sign in to view the complete answer paper.

Part B

210 marks
a

Let f(x) = 3x^2, when 0 ≤ x ≤ 1 be the probability density function of a continuous random variable X. Determine a and b such that i) P(X≤a) = P(X>a) ii) P(X>b) = 0.05.

5 marks

Answer

Sign in to view the complete answer paper.

b

Probability density function of a random variable X is

f(x) = (1/2)sin(x), 0 ≤ x ≤ π; 0, elsewhere
. Find the mean, mode and median of the distribution.

5 marks

Answer

Sign in to view the complete answer paper.

OR
310 marks
a

A die is cast until 6 appears. What is the probability that it must be cast more than 5 times?

5 marks

Answer

Sign in to view the complete answer paper.

b

The marks obtained in mathematics by 1000 students is normally distributed with mean 78% and standard deviation 11%. Determine: i) What was the highest mark obtained by the lowest 25% students? ii) Within what limit did the middle 90% of the student lie?

5 marks

Answer

Sign in to view the complete answer paper.

410 marks
a

Explain why the larger variance is placed in the numerator of the statistic F. Discuss the application of F-test in testing if two variances are homogenous.

5 marks

Answer

Sign in to view the complete answer paper.

b

A sample of 11 rats from a central population had an average blood viscosity of 3.92 with a standard deviation of 0.61. Estimate the 95% confidence limits for the mean blood viscosity of the population.

5 marks

Answer

Sign in to view the complete answer paper.

OR
510 marks
a

The mean voltage of a battery is 15 and standard deviation 0.2. Find the probability that four such batteries connected in series will have a combined voltage of 60.8 or more volts.

5 marks

Answer

Sign in to view the complete answer paper.

b

Discuss critical region and level of significance with example.

5 marks

Answer

Sign in to view the complete answer paper.

610 marks
a

Suppose the diameter of motor shafts in a lot have a mean of 0.249 inches and standard deviation if 0.003 inches. The inner diameter of bearings in another lot have a mean of 0.255 inches and standard deviation of 0.002 inches. If a shaft and bearing are selected at random, find the probability that the shaft will not fit inside the bearing. Assume that both dimensions are normally distributed.

5 marks

Answer

Sign in to view the complete answer paper.

b

A sample of 400 items is taken from a normal population whose mean is 4 and variance 4. If the sample mean is 4.45, can the samples be regarded as a simple sample?

5 marks

Answer

Sign in to view the complete answer paper.

OR
7

In a sample of 600 students of a certain college 400 are found to use ball pens. In another college from a sample of 900 students 450 were found to use ball pens. Test whether two colleges are significantly different with respect to the habit of using ball pens?

10 marks

Answer

Sign in to view the complete answer paper.

8

Estimate y at x = 5 by fitting a least squares curve of the form

y = b / [x(x-a)]
to the following data

x3.64.86.07.28.49.610.8
y0.830.310.170.100.070.050.04
10 marks

Answer

Sign in to view the complete answer paper.

OR
9

Show that the Gauss-Seidel methods diverge for solving the system of equations 2x + 3y + z = -1; 3x + 2y + 2z = 1; x + 2y + 2z = 6.

10 marks

Answer

Sign in to view the complete answer paper.

10

Find the successive approximate solution of the differential equation y' = y, y(0) = 1 by Picard’s method and compare it with exact solution.

10 marks

Answer

Sign in to view the complete answer paper.

OR
11

Use Runge-Kutta method of order four to find y when x = 0.6 in steps of 0.2 given that

dy/dx = 1 + y^2, y(0) = 0
.

10 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the question paper of MATHEMATICS - III from JNTU for B.Tech ME I Year I-II, August - 2018, R16.

For the solved question paper with answers, switch to the Answer Paper tab.

Other papers in MATHEMATICS - III

Loading Question Papers

Fetching data from database...